OpenStudy (anonymous):

PLEASE HELP! I'm failing Math!!! Axis of symmetry... g(x) = 5(x – 2)2 + 4 Identify the axis of symmetry for the graph of the following function. Enter your answer like this example: x = 1. g(x) = 5(x – 2)2 + 4 I don't know how to do this, and I am horrible with math.

Nnesha (nnesha):

\[\large\rm y=a(x-h)^2+k\] vertex form equation of parabola where (h,k) is the vertex h represent x-coordinate and k represent y-coordinate of the vertex x-coordinate = h= axis of symmetry

OpenStudy (anonymous):

Unfortunately... It sounds like Chinese to me. I have no idea what that means! D:

OpenStudy (campbell_st):

well basically the vertex is at (h, k) the line of symmetry for this parabola is at x = h so compare your equation the the parabola described by @Nnesha and identify the value of h,

Nnesha (nnesha):

sorry if math question sounds like Chinese to you \o^_^o/ \[\large\rm g(x) = 5(x – 2)^2 + 4\] this quadratic equation is written in vertex form of parabola \[\rm y=a(x-h)^2+k\]